x^1/3 multiplied by x^5/2, divided by x^-1/5
-- the directions on the problem say "simplify the expression. Express the answer so that only positive exponents occur. Assume that all variables are positive when they appear."
Thank you! Kim

Answer by sdmmadam@yahoo.com(530)

x^1/3 multiplied by x^5/2, divided by x^-1/5
-- the directions on the problem say "simplify the expression. Express the answer so that only positive exponents occur. Assume that all variables are positive when they appear."
x^1/3 multiplied by x^5/2, divided by x^-1/5
(x^1/3 X x^5/2)/( x^-1/5)
=x^(1/3+5/2)X [1/x^(-1/5)] { [something]/t = {that something X(1/t) }(example: (a/b) = aX(1/b) )
=x^(1/3+5/2)X x^(1/5) (using 1/[a^(-m)] =(a)^(m)
=x^[(2X1+3X5)/6] X [x^(1/5)] [using (a)^m X(a)^n = (a)^(m+n)] (and finding the lcm)
=x^(17/6) X x^(1/5)
=x^(17/6+1/5) [using (a)^m X(a)^n = (a)^(m+n)]
=x^[(5X17+6X1)/30] (finding the lcm)
=x[(85+6)/30]
=x^(91/30)